# On the Impact of Uncertainty and Calibration on Likelihood-Ratio Membership Inference Attacks

### v4

@@ V-C Evaluating MIA with DS Disclosure
Further results are described in Appendix -G.
@@ VI Conclusions
Figure 12: True advantage (14) and bounds (22) for the attacker’s advantage with CV, TLC, and DS observations in high TNR regime, $\alpha=0.999$, as a function of calibration error, $\Delta$; of epistemic uncertainty, $\epsilon_{e}$; and of aleatoric uncertainty, $\epsilon_{a}$.
@@
Accordingly, we evaluate the attack performance $\mathrm{Adv}_{0.999}=0.999-\beta_{0.999}$ at high TNR ($\alpha=0.999$), using the same settings as in Sec. V-B and Sec. V-C. Fig. 12 shows the attacker’s advantage with CV, TLC, and DS observations when $\alpha=0.999$ as a function of calibration error ($\Delta$), epistemic uncertainty ($\epsilon_{e}$), and aleatoric uncertainty ($\epsilon_{a}$). Fig. 13 illustrates the prediction set size and attacker’s advantage for DS disclosure when $\alpha=0.999$ as a function of the confidence threshold ($q$).
@@
Figure 14: The impact of the number of classes, of the subgroup size, and of the subgroup distribution on model calibration.
@@
We evaluate the relative calibration error as defined in (24) of our paper. Specifically, the expected confidence $\mathbb{E}_{f^{\mathrm{in}}}\left[p_{0}\right]$ is calculated by averaging the model’s predicted probabilities on training data with uncorrupted labels. As shown in Fig. 14 (top), the relative calibration error $\Delta$ increases with the number of classes. Our theoretical analysis, consistent with Fig. 7 in our paper, shows that a higher calibration error leads to a more pronounced MIA advantage. This explains why increasing class divisions improves MIA performance: it raises the model’s calibration error, aligning with empirical findings in [1]. With a fixed training data set size, more classes mean fewer samples per class, which, given sufficient model capacity and training resources, increases overfitting and calibration error. This, in turn, increases the vulnerability of the model to attacks, supporting our theoretical conclusions.
@@
To assess the impact of subgroup distribution on MIA performance, we design an experiment with fixed class sizes of $400$ samples per class on the MNIST data set. For class 00, we induce controlled distribution shifts using affine transformations, including rotation, translation, scaling, and shearing. The transformation intensity is adjusted by a scaling factor $\lambda\in\left\{0.1,0.2,0.3,0.4,0.5\right\}$, which linearly determines the magnitude of each transformation: rotation up to ±$\pm 40\lambda^{\circ}$, translation by ±$\textpm 20\lambda\%$ of the image size, scaling from $1.0$ to $1.0+0.5\lambda$, and shearing up to ±$\pm 20\lambda^{\circ}$. This controlled variation progressively increases the distributional divergence between the modified class 00 and unmodified classes $1-9$, allowing a quantitative analysis of how subgroup distribution influences MIA performance. As shown in Fig. 14 (bottom), the relative calibration error increases as the feature distribution of class 0 deviates further from the majority classes due to stronger affine transformations. The more distinct the subgroup’s distribution, the more the model memorizes its unique patterns, leading to overfitting and higher calibration error. This effect is more pronounced in smaller subgroups, where limited data prevents the model from generalizing effectively.

### v5

@@ V-C Evaluating MIA with DS Disclosure
Further experimental results are provided in Appendix -G, and discussions on performance under large-class settings are included in Appendix -H.
@@ VI Conclusions
Figure 12: True advantage (14) and bounds (22) for the attacker’s advantage with CV, TLC, and DS observations in high TNR regime, $\alpha=0.999$, as a function of calibration error, $\Delta$ (left); of aleatoric uncertainty, $\epsilon_{a}$ (middle); and of epistemic uncertainty, $\epsilon_{e}$ (right).
@@
Accordingly, we evaluate the attack performance $\mathrm{Adv}_{0.999}=0.999-\beta_{0.999}$ at high TNR ($\alpha=0.999$), using the same settings as in Sec. V-B and Sec. V-C. Fig. 12 shows the attacker’s advantage with CV, TLC, and DS observations when $\alpha=0.999$ as a function of calibration error ($\Delta$), aleatoric uncertainty ($\epsilon_{a}$), and epistemic uncertainty ($\epsilon_{e}$). Fig. 13 illustrates the prediction set size and attacker’s advantage for DS disclosure when $\alpha=0.999$ as a function of the confidence threshold ($q$).
@@
Figure 14: The impact of the number of classes (left), of the subgroup size (middle), and of the subgroup distribution on model calibration (right).
@@
We evaluate the relative calibration error as defined in (24) of our paper. Specifically, the expected confidence $\mathbb{E}_{f^{\mathrm{in}}}\left[p_{0}\right]$ is calculated by averaging the model’s predicted probabilities on training data with uncorrupted labels. As shown in Fig. 14 (left), the relative calibration error $\Delta$ increases with the number of classes. Our theoretical analysis, consistent with Fig. 7 in our paper, shows that a higher calibration error leads to a more pronounced MIA advantage. This explains why increasing class divisions improves MIA performance: it raises the model’s calibration error, aligning with empirical findings in [1]. With a fixed training data set size, more classes mean fewer samples per class, which, given sufficient model capacity and training resources, increases overfitting and calibration error. This, in turn, increases the vulnerability of the model to attacks, supporting our theoretical conclusions.
@@
To assess the impact of subgroup distribution on MIA performance, we design an experiment with fixed class sizes of $400$ samples per class on the MNIST data set. For class 00, we induce controlled distribution shifts using affine transformations, including rotation, translation, scaling, and shearing. The transformation intensity is adjusted by a scaling factor $\lambda\in\left\{0.1,0.2,0.3,0.4,0.5\right\}$, which linearly determines the magnitude of each transformation: rotation up to ±$\pm 40\lambda^{\circ}$, translation by ±$\textpm 20\lambda\%$ of the image size, scaling from $1.0$ to $1.0+0.5\lambda$, and shearing up to ±$\pm 20\lambda^{\circ}$. This controlled variation progressively increases the distributional divergence between the modified class 00 and unmodified classes $1-9$, allowing a quantitative analysis of how subgroup distribution influences MIA performance. As shown in Fig. 14 (right), the relative calibration error increases as the feature distribution of class 0 deviates further from the majority classes due to stronger affine transformations. The more distinct the subgroup’s distribution, the more the model memorizes its unique patterns, leading to overfitting and higher calibration error. This effect is more pronounced in smaller subgroups, where limited data prevents the model from generalizing effectively.
@@
-H Performance under Large-Class Settings
Figure 15: Empirical distributions of the confidence level $p_{0}$ and corresponding marginal of the fitted Dirichlet distributions for models trained without (left) and with (right) the target sample $(x_{0},y_{0})$ for CIFAR-100 data set with ResNet-50.
Figure 16: True advantage (14) and bounds (22) for the attacker’s advantage with TLC disclosure as a function of calibration error, $\Delta$ (left, $\epsilon_{a}=0.55$, $\epsilon_{e}=0.3$, $q=0.2$); of aleatoric uncertainty, $\epsilon_{a}$ (middle, $\Delta=0.5$, $\epsilon_{e}=0.3$, $q=0.2$); and of epistemic uncertainty, $\epsilon_{e}$ (right, $\Delta=0.5$, $\epsilon_{a}=0.55$, $q=0.4$).
To demonstrate the applicability of our framework in large-scale settings for current LiRA-style attacks [2, 16, 17, 18], we conduct experiments under TLC disclosure on CIFAR-100. We randomly select $N^{\mathrm{tr}}=4000$ images for data sets $\mathcal{D}^{\mathrm{out}}$ and $\mathcal{D}^{\mathrm{in}}$, and fine-tune a pre-trained ResNet-50 using SGD with a learning rate of $0.001$, momentum of $0.9$, and weight decay of $0.0005$ for $100$ epochs. Following the methodology in Sec. V-A, we extract softmax outputs for class-00 samples and fit their distributions using the two-parameter Beta form in equation (33), where $\gamma_{0}^{z}$ and $\overline{\gamma}_{0}^{z}$ represent the parameters for the true and aggregated non-true labels. As shown in Fig. 15, the fitted distributions closely match the empirical curves, validating the flexibility of our modeling assumption for large $K$.
The results in Fig. 16 further confirm that under large $K$, the trends under the TLC setting remain consistent with those observed in Fig. 7–Fig. 9 for TLC. This further supports the applicability of our theoretical framework to current LiRA-style attacks [2, 16, 17, 18], even in settings with a large number of classes.